Suppose the average length of time to finish a homework problem is 50 minutes. If a random sample of 9 students' times has standard deviation 6, what is the approximate probability the average time of the sampled students to complete the homework will be more than 53.7 minutes?
Answer #1

Solution :

\sigma\bar x = \sigma / \sqrt n = 6 / \sqrt 9 = 2

P(\bar x > 53.7) = 1 - P(\bar x < 53.7)

= 1 - P[(\bar x - \mu \bar x ) / \sigma \bar x < (53.7 - 50) / 2]

= 1 - P(z < 1.85)

= 1 - 0.9678

= 0.0322

Probability = 0.0322